OpenStudy (anonymous):

assuming an angle in quadrant I , evaluate csc(cot^-1 4/3) a) 4/5 b)5/4 c)5/3 d)3/5

OpenStudy (anonymous):

@ajprincess could you please please help me!?

OpenStudy (anonymous):

They want \[ \csc(\cot^{-1}(4/3)) \]?

OpenStudy (anonymous):

Remember that \(\cot(\alpha)=\cos(\alpha)/\sin(\alpha)\)

OpenStudy (anonymous):

yes thats what they're asking

OpenStudy (anonymous):

how will that identity help?

OpenStudy (anonymous):

Because this: 

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We can see that 4 corresponds to sin, and 3 corresponds to cos.

OpenStudy (anonymous):



OpenStudy (anonymous):

So first you need to find the hypotenuse.

OpenStudy (anonymous):

Remember that cos is adjacent over hypotenuse. csc is just the reciprocal of cos, ie hypotenuse over adjacent. We just need the hypotenuse.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

is it 5/3 @wio ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

cot^-1=4/3 tan x=4/3 hypotnuse5 sinx=4/5 cosec=5/4

OpenStudy (anonymous):

so its 5/4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hope it helps